Main INDEX
Monthly INDEX
PREV
NEXT
User name Ulmer
Log entry time 16:07:39 on November 3,2002
Entry number 89513
keyword=Day Shift Summary
Start shift with beam.
08:10: MCC calls. They have beam trips related to vacuum problem in Hall A. May be a "phantom" problem, but they want us to check our systems: something between the shield wall and Hall A dump. I see nothing on Hall A vacuum screens worth mentioning, though am not really familiar with these screens.
At same time, notice that our BPMB X reading on General Tools screen is almost 1 mm off. Tell MCC we want beam centered; it appears they were unaware that this reading should be zero. I don't know how long we've been running with this uncentered beam. Now it's centered in both X and Y.
08:50: The 1C12 X strip chart shows an offscale reading (beyond +/-0.2mm) for the entire owl shift. At around 8am the same chart shows a stable value near 0. A call to MCC reveals that the fast feedback was turned ON around 8am. They do not know how long it was OFF. The strip chart indicates it's been off for at least 9 hours.
09:00: Upon Hall C request, 1/2 wave plate will be removed.
09:30: Lower current from 100 uA to 85 uA in a hope to reduce # of beam trips. As of 10:30 it appears that this had no effect. MCC says we might be able to get about a factor of two faster ramp up rate which, according to their estimate, would reduce the ramp percentage from about 25% to 12% (i.e. given the number of trips and the time of ramping for each, we spend about 25% of any hour just ramping up).
10:35: As of 10:20 we have had very little beam, so hold off on asking MCC for faster ramp up. Apparently, Paul from Hall C asked for 4 uA and they thought is was the Hall A Paul. Anyway, current is back to normal.
10:45: MCC will increase ramp-up rate from "3" to "4.5". We will monitor target for next couple of trips and then possibly increase to "6".
11:40: We see very little improvement from lower current. So we asked MCC to go to 100 uA. Also asked to bring ramp-up rate from "4.5" to "6". Now running at 100 uA. They have also made changes to path length correction (~3 deg).
12:45: MCC calls reporting trouble with RF zone - vacuum problem --> no beam.
End run 2936 at 12:55.
13:07: Beam back.
14:10: MCC is now increasing ramp-up rate to "6".
14:20: MCC asked to take energy lock away to test whether "fantom" vacuum trips are related to energy lock (they appeared to start when the lock was turned on this morning). They would like to test whether over 1/2 hr or so the # of these trips reduces. I say OK, since MCC will keep energy stable "by hand", but we are also watching the 1C12 X position strip chart.
14:35: We get energy lock back (see Halog # 89509).
Fairly steady beam (though several short trips) for remainder of shift.